PLEASE MAKE THE ANSWER CLEAR AND VISIBLE. I GIVE THUMBS UP

Lazare Corporation expects an EBIT of $17,100 every year forever. Lazare currently has no debt, and its cost of equity is 10%. The firm can borrow at 7%. (Do not round intermediate calculations. Round the final answers to 2 decimal places. Omit $ sign in your response.)

a. If the corporate tax rate is 35%, what is the value of the firm?

Value of the firm            $

b. What will the value be if the company converts to 50% debt?

Value of the firm            $

c. What will the value be if the company converts to 100% debt?

Value of the firm            $
Answer #1

a. The value of the firm is computed as shown below:

[ EBIT ( 1 - tax rate ) ] / cost of equity

= [ 17,100 ( 1 - 0.35 ) ] / 0.10

= 111,150

b. The value of the firm is computed as shown below:

= 111,150 + 111,150 x tax rate x 50%

= 111,150 + 111,150 x 0.35 x 0.50

= 130,601.25

c. The value of the firm is computed as shown below:

= 111,150 + 111,150 x tax rate x 1

= 111,150 + 111,150 x 0.35 x 1

= 150,052.5

Feel free to ask in case of any query relating to this question